# On the free metabelian Novikov and metabelian Lie-admissible algebras

### v2

@@ Abstract.
Key words and phrases:
Novikov algebra, Lie-admissible algebra, metabelian identity, free algebra, polynomial identities
2010 Mathematics Subject Classification:
17A30, 17A50, 17D25
1. Introduction
@@
2. Free metabelian Novikov algebra
@@
Let $X=\{x_{1},x_{2}\ldots\}$ be a countable set of generators. We denote by $\mathrm{Nov}\langle X\rangle$ and $\mathrm{M}\mathrm{Nov}\langle X\rangle$ the free Novikov and free metabelian Novikov algebra, respectively.
@@
2.1 Lemma.
The algebra $\mathrm{M}\mathrm{Nov}\langle X\rangle$ satisfies the following identities:
@@
2.2 Theorem.
The set $\bigcup\mathcal{N}_{i}$ is a linear basis of the algebra $\mathrm{M}\mathrm{Nov}\langle X\rangle$.
Proof.
The statement for degrees less than 5 can be verified by direct calculations. For $n\geq 5$, firstly, we show that every monomial of $\mathrm{M}\mathrm{Nov}\langle X\rangle$ can be written as a sum of monomials from the set $\bigcup_{i\geq 5}\mathcal{N}_{i}$. By Lemma (2.1), we obtain that every monomial except $(\cdots((x_{i_{1}}x_{i_{2}})x_{i_{3}})\cdots)x_{i_{n}}$ is equal to 00. By (2) and Lemma (2.1), we have
$(\cdots((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{m-1}})x_{i_{m}})\cdots)x_{n}=(\cdots((\cdots(x_{i_{1}}x_{i_{2}})\cdots)(x_{i_{m-1}}x_{i_{m}}))\cdots)x_{n}+\\ (\cdots(((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{m}})x_{i_{m-1}})\cdots)x_{n}-(\cdots((\cdots(x_{i_{1}}x_{i_{2}})\cdots)(x_{i_{m}}x_{i_{m-1}}))\cdots)x_{n}=\\ (\cdots(((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{m}})x_{i_{m-1}})\cdots)x_{n},$
i.e., the generators $x_{i_{2}}$, $x_{i_{3}},\ldots$, $x_{i_{n}}$ are rearrangeable. From these equations, we get that every monomial of $\mathrm{M}\mathrm{Nov}\langle X\rangle$ which has a degree greater than 4 can be written as a sum of $\bigcup_{i\geq 5}\mathcal{N}_{i}$.
@@
where $n>4$ and $i_{2}\leq\ldots\leq i_{k}\leq j\leq i_{k+1}\leq\ldots\leq i_{n}$. Up to degree 4, we define multiplication in $A\langle X\rangle$ that is consistent with identities (1), (2) and (3). By straightforward calculation, we can check that an algebra $A\langle X\rangle$ satisfies to (1), (2) and (3) identities. It remains to note that $A\langle X\rangle\cong\mathrm{M}\mathrm{Nov}\langle X\rangle$. ∎
3. Symmetric polynomials of the free metabelian Novikov algebra
@@
where $\sigma\in S_{n}$. Let us define a set of polynomials $\mathcal{P}$ in $\mathrm{M}\mathrm{Nov}\langle X\rangle$ as follows:
@@
3.1 Example.
For multilinear part of $\mathrm{M}\mathrm{Nov}\langle X\rangle$, we obtain
@@
Proof.
For $n=1,2$, the result is obvious. For $n\geq 3$, we use the fact that every Novikov algebra is embeddable into commutative algebra with derivation and spanning elements of Novikov algebra in commutative algebra with derivation are monomials of weight $-1$ [4, 16]. The weight function $\mathrm{wt}(u)\in\mathbb{Z}$ is defined on monomials of commutative algebra with derivation $D$ by induction as follows,
$\mathrm{wt}(x)=-1,\quad x\in X;$
$\mathrm{wt}(d(u))=\mathrm{wt}(u)+1;\quad\mathrm{wt}(uv)=\mathrm{wt}(u)+\mathrm{wt}(v).$
For simplicity, we denote $D(x)$ and $D^{n}(x)$ by $x^{\prime}$ and $(D^{n-1}(x))^{\prime}$, respectively. For the multilinear part to describe the space of symmetric polynomials of $\mathrm{Nov}\langle X\rangle$ in degree $3$, we need to find the space of the symmetric polynomials of the differential commutative algebra of weight $-1$ in degree $3$. This is the monomials of the form
@@
For degree $4$, we have $\mathrm{M}\mathrm{Nov}\langle X\rangle/\{(ab)(cd)\}\cong\mathrm{Com}\langle X\rangle^{(D)}_{-1}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$, where $\mathrm{Com}\langle X\rangle^{(D)}_{-1}$ is a space of differential commutative algebra of weight $-1$. Hence, in $\mathrm{Com}\langle X\rangle^{(D)}_{-1}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ we rewrite all monomials of the form $x_{i_{1}}^{\prime\prime}x_{i_{2}}^{\prime}x_{i_{3}}x_{i_{4}}$ to $-x_{i_{1}}^{\prime}x_{i_{2}}^{\prime}x_{i_{3}}^{\prime}x_{i_{4}}$. It remains to note that
@@
which gives that we rewrite monomial $a^{\prime}b^{\prime}d^{\prime}c$ to $a^{\prime}b^{\prime}c^{\prime}d$. Finally, the remained monomials of weight $-1$ are $x_{i_{1}}x_{i_{2}}x_{i_{3}}x_{i_{4}}^{\prime\prime\prime}$ and $x_{j_{1}}x_{j_{2}}^{\prime}x_{j_{3}}^{\prime}x_{j_{4}}^{\prime}$, where $i_{1}\leq i_{2}\leq i_{3}$ and $j_{1}\leq j_{2}\leq j_{3}\leq j_{4}$. By Theorem 2.2, these monomials are linearly independent, and for multilinear part symmetric polynomials of $\mathrm{Com}\langle X\rangle^{(D)}_{-1}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ are
@@
By Theorem 2.2, starting from degree $5$, all monomials of $\mathrm{Com}\langle X\rangle^{(D)}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ of weight $-1$ except $x_{i}^{(n-1)}x_{j_{n-1}}\ldots x_{j_{1}}$ are equal to 00, where $x_{i}^{(n-1)}=(x_{i}^{(n-2)})^{\prime}$. Hence, starting from degree $5$, symmetric polynomials of $\mathrm{Com}\langle X\rangle^{(D)}/$$\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ of weight $-1$ are
@@
which correspond to $p_{n}$. ∎
4. Free metabelian Lie-admissible algebra
@@
Calculating the dimension of operad $\mathrm{M}\mathrm{Lie}$ by means of the package [7], we get the following result:
@@
$\dim(\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}(n))$ 1 2 11 77 679 7184 88668
@@
$\dim(\mathrm{Lie}\textrm{-}\mathrm{adm}(n))$ 1 2 11 101 1299 21484 434314
@@
5. Symmetric polynomials of the free metabelian Lie-admissible algebra
@@
The sum of such monomials gives polynomial $p_{(*,*,\ldots,*,n)}$. For $n=1,2$, we set $p_{(1)}=x_{1}+x_{2}+\ldots+x_{n}$ and $p_{(2)}=\sum_{i\neq j}x_{i}\circ x_{j}$.
5.2 Example.
For multilinear part of $\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}\langle X\rangle$, let us construct $(\bullet_{x_{i_{1}}},\circ_{x_{i_{2}}},$ $x_{i_{3}}{)}_{(\bullet,\circ)}$, ${(\circ_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})}_{(\bullet,\circ)}$ and ${(\circ_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})}_{(\bullet,\circ)}$.
@@
$\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}_{\geq 3}\langle X\rangle=\oplus_{(*,*,\ldots,*)}(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$
as a vector space, where $\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}_{\geq 3}\langle X\rangle$ is a multilinear part of the free metabelian Lie-admissible algebra of degree greater than $2$. For example, if $n=3$ then
$\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}_{3}\langle X\rangle=(\bullet_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})_{(\bullet,\bullet)}\oplus(\bullet_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})_{(\bullet,\circ)}\oplus\\ (\circ_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})_{(\circ,\bullet)}\oplus(\circ_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})_{(\circ,\circ)}.$
Moreover, each monomial of the space $(*_{x_{i_{1}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$ under action $S_{n}$ belongs to the same space. It remains to note that $p_{(*,*,\ldots,*,n)}$ is a symmetric polynomial, and for each space $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},$ $x_{i_{n}})_{(*,*,\ldots,*)}$ there is one unique symmetric polynomial which is $p_{(*,*,\ldots,*,n)}$. ∎
@@ ACKNOWLEDGMENTS
This work was supported by UAEU grant G00004614.

### v3

@@ Abstract
keywords:
Novikov algebra, Lie-admissible algebra, metabelian identity, free algebra, polynomial identities.
[…]
1 Introduction
@@
2 Free metabelian Novikov algebra
@@
Let $X=\{x_{1},x_{2}\ldots\}$ be a countable set of generators. We denote by $\operatorname{{Nov}}\langle X\rangle$ and $\operatorname{{M}}\operatorname{{Nov}}\langle X\rangle$ the free Novikov and free metabelian Novikov algebra, respectively.
@@
Lemma 2.1.
The algebra $\operatorname{{M}}\operatorname{{Nov}}\langle X\rangle$ satisfies the following identities:
@@
Theorem 2.3.
The set $\bigcup\mathcal{N}_{i}$ is a linear basis of the algebra $\operatorname{{M}}\operatorname{{Nov}}\langle X\rangle$.
Proof 2.4.
The statement for degrees less than 5 can be verified by direct calculations. For $n\geq 5$, firstly, we show that every monomial of $\operatorname{{M}}\operatorname{{Nov}}\langle X\rangle$ can be written as a sum of monomials from the set $\bigcup_{i\geq 5}\mathcal{N}_{i}$. By Lemma 2.1, we obtain that every monomial except $(\cdots((x_{i_{1}}x_{i_{2}})x_{i_{3}})\cdots)x_{i_{n}}$ is equal to 00. By (2) and Lemma 2.1, we have
$(\cdots((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{m-1}})x_{i_{m}})\cdots)x_{i_{n}}=(\cdots((\cdots(x_{i_{1}}x_{i_{2}})\cdots)(x_{i_{m-1}}x_{i_{m}}))\cdots)x_{i_{n}}+\\ (\cdots(((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{m}})x_{i_{m-1}})\cdots)x_{i_{n}}-(\cdots((\cdots(x_{i_{1}}x_{i_{2}})\cdots)(x_{i_{m}}x_{i_{m-1}}))\cdots)x_{i_{n}}=\\ (\cdots(((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{m}})x_{i_{m-1}})\cdots)x_{i_{n}},$
i.e., the generators $x_{i_{2}}$, $x_{i_{3}},\ldots$, $x_{i_{n}}$ are rearrangeable. From these equations, we get that every monomial of $\operatorname{{M}}\operatorname{{Nov}}\langle X\rangle$ which has a degree greater than 4 can be written as a sum of $\bigcup_{i\geq 5}\mathcal{N}_{i}$.
@@
where $n>4$ and $i_{2}\leq\ldots\leq i_{k}\leq j\leq i_{k+1}\leq\ldots\leq i_{n}$. Up to degree 4, we define multiplication in $A\langle X\rangle$ that is consistent with identities (1), (2) and (3). By straightforward calculation, we can check that an algebra $A\langle X\rangle$ satisfies to (1), (2) and (3) identities. It remains to note that $A\langle X\rangle\cong\operatorname{{M}}\operatorname{{Nov}}\langle X\rangle$.
3 Symmetric polynomials of the free metabelian Novikov algebra
@@
where $\sigma\in S_{n}$. Let us define a set of polynomials $\mathcal{P}$ in $\operatorname{{M}}\operatorname{{Nov}}\langle X\rangle$ as follows:
@@
Example 3.1.
For multilinear part of $\operatorname{{M}}\operatorname{{Nov}}\langle X\rangle$, we obtain
@@
Proof 3.3.
For $n=1,2$, the result is obvious. For $n\geq 3$, we use the fact that every Novikov algebra is embeddable into commutative algebra with derivation and spanning elements of Novikov algebra in commutative algebra with derivation are monomials of weight $-1$ [4, 16]. The weight function $\operatorname{{wt}}(u)\in\mathbb{Z}$ is defined on monomials of commutative algebra with derivation $D$ by induction as follows,
$\operatorname{{wt}}(x)=-1,\quad x\in X;$
$\operatorname{{wt}}(d(u))=\operatorname{{wt}}(u)+1;\quad\operatorname{{wt}}(uv)=\operatorname{{wt}}(u)+\operatorname{{wt}}(v).$
For simplicity, we denote $D(x)$ and $D^{n}(x)$ by $x^{\prime}$ and $(D^{n-1}(x))^{\prime}$, respectively. For the multilinear part to describe the space of symmetric polynomials of $\operatorname{{Nov}}\langle X\rangle$ in degree $3$, we need to find the space of the symmetric polynomials of the differential commutative algebra of weight $-1$ in degree $3$. This is the monomials of the form
@@
For degree $4$, we have
$\operatorname{{M}}\operatorname{{Nov}}\langle X\rangle/\{(ab)(cd)\}\cong\operatorname{Com}\langle X\rangle^{(D)}_{-1}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\},$
where $\operatorname{Com}\langle X\rangle^{(D)}_{-1}$ is a space of differential commutative algebra of weight $-1$. Hence, in degree $4$ of $\operatorname{Com}\langle X\rangle^{(D)}_{-1}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ we rewrite all monomials of the form $x_{i_{1}}^{\prime\prime}x_{i_{2}}^{\prime}x_{i_{3}}x_{i_{4}}$ to $-x_{i_{1}}^{\prime}x_{i_{2}}^{\prime}x_{i_{3}}^{\prime}x_{i_{4}}$. It remains to note that
@@
which gives that we rewrite monomial $a^{\prime}b^{\prime}d^{\prime}c$ to $a^{\prime}b^{\prime}c^{\prime}d$. Finally, the remained monomials of weight $-1$ are $x_{i_{1}}x_{i_{2}}x_{i_{3}}x_{i_{4}}^{\prime\prime\prime}$ and $x_{j_{1}}x_{j_{2}}^{\prime}x_{j_{3}}^{\prime}x_{j_{4}}^{\prime}$, where $i_{1}\leq i_{2}\leq i_{3}$ and $j_{1}\leq j_{2}\leq j_{3}\leq j_{4}$. By Theorem 2.3, these monomials are linearly independent, and for multilinear part symmetric polynomials of $\operatorname{Com}\langle X\rangle^{(D)}_{-1}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ are
@@
By Theorem 2.3, starting from degree $5$, all monomials of $\operatorname{Com}\langle X\rangle^{(D)}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ of weight $-1$ except $x_{i}^{(n-1)}x_{j_{n-1}}\ldots x_{j_{1}}$ are equal to 00, where $x_{i}^{(n-1)}=(x_{i}^{(n-2)})^{\prime}$. Hence, starting from degree $5$, symmetric polynomials of $\operatorname{Com}\langle X\rangle^{(D)}/$$\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ of weight $-1$ are
@@
which correspond to $p_{n}$.
4 Free metabelian Lie-admissible algebra
@@
Calculating the dimension of operad $\operatorname{{M}}\operatorname{Lie}$ by means of the package [6], we get the following result:
@@
$\dim(\operatorname{{M}}\operatorname{Lie}\textrm{-}\operatorname{adm}(n))$ 1 2 11 77 679 7184 88668
@@
$\dim(\operatorname{Lie}\textrm{-}\operatorname{adm}(n))$ 1 2 11 101 1299 21484 434314
@@
5 Symmetric polynomials of the free metabelian Lie-admissible algebra
@@
The sum of such monomials gives polynomial $p_{(*,*,\ldots,*,n)}$. For $n=1,2$, we set
$p_{(1)}=x_{1}+x_{2}+\ldots+x_{n}\quad\textup{ and }\quad p_{(2)}=\sum_{i\neq j}x_{i}\circ x_{j}.$
Example 5.2.
For multilinear part of $\operatorname{{M}}\operatorname{Lie}\textrm{-}\operatorname{adm}\langle X\rangle$, let us construct $(\bullet_{x_{i_{1}}},\circ_{x_{i_{2}}},$ $x_{i_{3}}{)}_{(\bullet,\circ)}$, ${(\circ_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})}_{(\bullet,\circ)}$ and ${(\circ_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})}_{(\bullet,\circ)}$.
@@
$\operatorname{{M}}\operatorname{Lie}\textrm{-}\operatorname{adm}_{\geq 3}\langle X\rangle=\oplus_{(*,*,\ldots,*)}(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$
as a vector space, where $\operatorname{{M}}\operatorname{Lie}\textrm{-}\operatorname{adm}_{\geq 3}\langle X\rangle$ is a multilinear part of the free metabelian Lie-admissible algebra of degree greater than $2$. For example, if $n=3$ then
$\operatorname{{M}}\operatorname{Lie}\textrm{-}\operatorname{adm}_{3}\langle X\rangle=(\bullet_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})_{(\bullet,\bullet)}\oplus(\bullet_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})_{(\bullet,\circ)}\oplus\\ (\circ_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})_{(\circ,\bullet)}\oplus(\circ_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})_{(\circ,\circ)}.$
Moreover, each monomial of the space $(*_{x_{i_{1}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$ under action $S_{n}$ belongs to the same space. It remains to note that $p_{(*,*,\ldots,*,n)}$ is a symmetric polynomial, and for each space $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},$ $x_{i_{n}})_{(*,*,\ldots,*)}$ there is one unique symmetric polynomial which is $p_{(*,*,\ldots,*,n)}$.
@@ Acknowledgments
This work was supported by UAEU grant G00004614. The authors are grateful to the anonymous referees for their valuable remarks that improved the text.